# FLrce: Resource-Efficient Federated Learning with Early-Stopping Strategy

### v2

@@ Abstract
Federated learning (FL) achieves great popularity in the Internet of Things (IoT) as a powerful interface to offer intelligent services to customers while maintaining data privacy. Under the orchestration of a server, edge devices (also called clients in FL) collaboratively train a global deep-learning model without sharing any local data. Nevertheless, the unequal training contributions among clients have made FL vulnerable, as clients with heavily biased datasets can easily compromise FL by sending malicious or heavily biased parameter updates. Furthermore, the resource shortage issue of edge devices also becomes a bottleneck. Due to overwhelming computation overheads generated by training deep-learning models on edge devices, and significant communication overheads for transmitting deep-learning models across the network, enormous amounts of resources are consumed in the FL process. This encompasses computation resources like energy and communication resources like bandwidth. To comprehensively address these challenges, in this paper, we present FLrce, an efficient FL framework with a relationship-based client selection and early-stopping strategy. FLrce accelerates the FL process by selecting clients with more significant effects, enabling the global model to converge to a high accuracy in fewer rounds. FLrce also leverages an early stopping mechanism that terminates FL in advance to save communication and computation resources. Experiment results show that, compared with existing efficient FL frameworks, FLrce improves the computation and communication efficiency by at least 47% and 43% respectively.
@@ 1 Introduction
Even though FL remarkably upholds data privacy, when it comes to application, the effectiveness of FL is usually compromised by the issues of resource shortage and client heterogeneity in real-world IoT networks. On one hand, due to the resource constraints of edge devices, the communication and computation resources of an IoT network are usually finite, presenting a critical challenge to address the resource shortage issue of FL in IoT settings [10, 11, 2]. As shown in Figure 1, first, an IoT network with limited bandwidth support can hardly withstand the exorbitant communication cost in FL when deep-learning models with massive amounts of parameters are communicated frequently between clients and the server [12, 2, 13]. Second, due to the significant energy consumption of training deep learning models on edge devices [14], FL usually generates high computation costs which can be unaffordable for IoT systems with deficient energy supply. To solve the resource shortage problem, an efficient FL framework that can reduce the consumption and communication costs without degrading accuracy is expected.
@@
Motivated by the limitation of the state-of-the-art, in this paper, we present FLrce, i.e. Federated Learning with relationship-based client selection and early stopping strategy, to simultanesly address the client heterogeneity and resource shortage problems of FL in IoT settings.
@@
  • We implement FLrce on a physical platform and evaluate FLrce on real-world datasets. The experiment results show that, compared with existing efficient FL frameworks, FLrce at least increases the accuracy by 3.9% and the efficiency of the computation/communication resource utilization by 47% and 43% respectively on average.
@@ 2.1 Federated Learning
Figure 3: Existing efficient FL methods (the blue bars) with a random client selection strategy easily become vulnerable to client heterogeneity.
Figure 4: Existing efficient FL methods consume more communication and computation resources to reach the same target accuracy.
@@ 2.2 Degraded Performance Caused by Client Heterogeneity
Due to the non-iid local data distribution of FL clients, the performance of efficient FL frameworks can be easily compromised by unbalanced training contributions from the heterogeneous clients. For illustration, we implement Fedcom [16], Fedprox [21], Dropout [25] and FLrce for the voice assisting task on the Google Speech dataset [34]. Unlike FLrce, Fedcom, Fedprox and Dropout all apply a random client selection strategy. To create a heterogeneous environment with unbalanced client contributions, data are unevenly distributed to clients following a Dirichlet distribution with parameter 0.1 [35]. We simply run each method for 20 rounds, and Figure 3 lists the accuracy of all methods. From Figure 3 we can see that existing efficient FL methods are significantly impeded by client heterogeneity owning to a random client selection strategy, as they achieve lower accuracy than FLrce, which applies a relationship-based client selection strategy.
@@ 2.3 Additional Resource Consumption Resulting from Low Learning Efficiency
Owing to the degraded performance caused by client heterogeneity, efficient FL methods inevitably require additional consumption of communication and computation resources to reach the same target accuracy. For illustration, we estimate the computation/communication costs of Fedcom, Fedprox, Dropout and FLrce to achieve the same target accuracy on the Google Speech dataset. From Figure 4 we can see that, existing efficient FL methods obtain much lower learning efficiency due to the unbalanced training contributions from heterogeneous clients. Specifically, Fedcom, Fedprox, and Dropout are expected to consume 1.35$\times$, 2.04$\times$, 5.88$\times$ the computation resource (e.g. energy [14, 10]) and 1.63$\times$, 2.43$\times$, 3.57$\times$ the communication resource (e.g. bandwidth [13]) to reach the same accuracy compared with FLrce.
@@ 3 Design of FLrce
In this section we introduce the proposed FLrce framework. Specifically, subsection 3.1 lays out a high-level demonstration of how FLrce works. Subsections 3.2 and 3.3 respectively explain the client selection strategy and the early stopping mechanism of FLrce. Subsection 3.4 summarizes all the details and presents a comprehensive overview of FLrce.
@@ 3.1 Solution Overview
A high-level workflow of FLrce is presented in Figure 5. 1⃝: At the beginning of round $t$, the server decides the set of active clients based on a client selection strategy $h$. $h$ takes $H$ as input and returns the server a set of clients $\mathbb{C}_{t}$, where $H$ is the heuristic map that stores the heuristic values of clients (see Algorithm 2). 2⃝: The server broadcasts the current global model $w^{t}$ to all selected clients. 3⃝: In parallel, each participating client $C_{k}$ ($C_{k}\in\mathbb{C}_{t}$) updates the received parameter by applying stochastic gradient descent (SGD) over its personal dataset (see Equation (3)). 4⃝: Each $C_{k}$ uploads the parameter updates $u^{t}_{k}$ to the server. 5⃝: For each $C_{k}\in\mathbb{C}_{t}$, the server writes the corresponding $u^{t}_{k}$ to $V_{k}$, and writes $t$ to $R_{k}$. $V$ is an update map that stores each client’s latest update, $R$ is a time map that stores each client’s last active round. 6⃝: The server aggregates the received updates and updates the global model parameter (see Equation (4)). 7⃝: The server evaluates the relationship between each selected $C_{k}$ and other clients, and writes the relationships into a relationship map $\Omega\in\mathbb{R}^{M\times M}$, using a heuristic function $g$(see Algorithm 1). 8⃝: The server computes each $C_{k}$’s latest heuristic value based on the newest $\Omega$ and updates the heuristic map $H$ (see Equation (7)). 9⃝: The server examines whether the early-stopping mechanism has been triggered. If so, the server terminates the FL process (see subsection 3.3). If not, the server moves ahead to round $t+1$ and restarts from step 1⃝. The summary of all mathematical notations used in this paper is listed in Table I.
@@ 3.2 Relationship-based Client Selection
$d_{p}>d_{o}$ indicates that the global model is possibly moving away from $C_{q}$’s local optimum after incorporating the update from $C_{p}$, then a negative relationship degree will be assigned between $C_{p}$ and $C_{q}$. If $d_{p}<d_{o}$, $C_{p}$ and $C_{q}$ will have a positive relationship degree for a similar reason. To control the relationship degree on the same scale as cosine similarity, we make the result in Equation (6) be no less than -1. The computation of Equation (6) will not infringe on data privacy, as it only requires clients’ parameter updates and the global model parameter.
Algorithm 1 Heuristic function $g$
1:update $u^{t}_{k}$, old map $\Omega$, update map $V$, time map $R$, round $t$, client index $k$
@@
6:         synchronous relationship modelling:
@@
9:         asynchronous relationship modelling:
@@
13:return new map $\Omega$
With the synchronous and asynchronous RM combined, the heuristic function $g$ is implemented in Algorithm 1.
Heuristic Values Computation. The heuristic value of a client $C_{k}$ reflects the importance of $C_{k}$ to the collaborative learning process. FLrce identifies the important clients as the ones that maintain more positive relationships with others. Therefore, we define the heuristic of $C_{k}$ as a summation of its relationship degrees with other clients:
$H_{k}=\sum_{j=1,j\neq k}^{M}\Omega_{k,j}$ (7)
@@
3.3 Early Stopping Mechanism
FLrce presents a novel early stopping mechanism settings that terminates FL ahead of the maximum training round. Unlike [33] that uses early stopping to prevent overfitting in centralized machine learning, the early stopping mechanism in FLrce aims to reduce the overall communication and computation resource consumption in decentralized FL settings. The point is that, as FL goes on, the global model tends to ”oscillate” between the important clients’ local optima. In this case, it becomes much more difficult for the global model to reach the global optimum, which triggers the ES mechanism to avoid further consumption of communication and computation resources.
For illustration, Figure 9 shows a simple two-dimensional example. Suppose a network with five clients $C_{1}\sim C_{5}$. $C_{2}$ and $C_{3}$ are the two most important clients as their local optima are closer to the global optimum than others. Given $P=2$, the server tends to to select $C_{2}$ and $C_{3}$ in an exploit round. Suppose two exploit rounds $t_{1}$ and $t_{2}$. At $t_{1}$, the global model $w^{t_{1}}$ is far away from the optimal model $w^{*}$, and the server finds $C_{2}$ and $C_{3}$ not conflicting as $cossim(u_{2}^{t_{1}},u_{3}^{t_{1}})>0$. On the contrary, $w^{t_{2}}$ is closer to $w^{*}$ and the server finds $C_{2}$ and $C_{3}$ conflicting as $cossim(u_{2}^{t_{2}},u_{3}^{t_{2}})<0$. Accordingly, we hypothesize that, when the most important clients conflict with each other, the global model is likely to be in close proximity to the optimal solution.
@@ 3.3 Early Stopping Mechanism
max iterations $T$, number of participants $P$, clients $\mathbb{C}$, explore-exploit factor $\phi$, ES threshold $\psi$
Globally, server $S$ executes:
1:initialize $w^{0}$, $H$, $\Omega$, $R$, $V$
2:for $t=1,2,...,T$ do
3:     client selection:
4:           $\mathbb{C}_{t}\leftarrow h(H,T,P,\mathbb{C},\phi)$triangleright Alg. 2
5:     for $k\in\{1,...,M\},C_{k}\in\mathbb{C}_{t}$ do
6:         send $w^{t}$ to $C_{k}$
7:         receive $u^{t}_{k}$ from $C_{k}$
8:         write update/round to the update/time map:
9:               $V_{k}\leftarrow u^{t}_{k},R_{k}\leftarrow t$
10:     end for
11:     aggregation:
12:           $w^{t+1}\leftarrow w^{t}+\sum_{k:C_{k}\in\mathbb{C}_{t}}p_{k}u_{k}^{t}$triangleright Eq. (4)
13:     for $k\in\{1,...,M\},C_{k}\in\mathbb{C}_{t}$ do
14:         update relationship:
15:               $\Omega\leftarrow g(u^{t}_{k},V,R,t,k,\Omega)$triangleright Alg. 1
16:         update heuristic:
17:               $H_{k}=\sum_{j=1,j\neq k}^{M}\Omega_{k,j}$triangleright Eq. (7)
18:     end for
19:     check early stopping:
20:     if $ES(t,\mathbb{C}_{t},P,\psi)==$ True then $\triangleright$ Alg. 3
21:         break
22:     end if
23:end for
24:return $w^{t+1}$
@@
(a) EMNIST
(b) Google Speech
(c) CIFAR10
(d) CIFAR100
Figure 10: Mean validation accuracy per round
Dataset Fedcom Fedprox Dropout FLrce
Accuracy Round Accuracy Round Accuracy Round Accuracy Round
EMNIST 78.2% 100 65.9% 100 8.5% 100 76.1% 44
Google Speech 55.8% 100 47.1% 100 21.6% 100 58.6% 52
CIFAR10 42.9% 100 39.5% 100 14.9% 100 53.8% 63
CIFAR100 32.3% 100 29.9% 100 13.7% 100 36.6% 47
Average 52.3% 100 45.6% 100 14.6% 100 56.2 % 52
TABLE III: Total rounds of training and final validation accuracy
3.4 Put All the Pieces Together

### v4

@@ Abstract
Federated Learning (FL) achieves great popularity in the Internet of Things (IoT) as a powerful interface to offer intelligent services to customers while maintaining data privacy. Under the orchestration of a server, edge devices (also called clients in FL) collaboratively train a global deep-learning model without sharing any local data. Nevertheless, the unequal training contributions among clients have made FL vulnerable, as clients with heavily biased datasets can easily compromise FL by sending malicious or heavily biased parameter updates. Furthermore, the resource shortage issue of the network also becomes a bottleneck. Due to overwhelming computation overheads generated by training deep-learning models on edge devices, and significant communication overheads for transmitting deep-learning models across the network, enormous amounts of resources are consumed in the FL process. This encompasses computation resources like energy and communication resources like bandwidth. To comprehensively address these challenges, in this paper, we present FLrce, an efficient FL framework with a relationship-based client selection and early-stopping strategy. FLrce accelerates the FL process by selecting clients with more significant effects, enabling the global model to converge to a high accuracy in fewer rounds. FLrce also leverages an early stopping mechanism that terminates FL in advance to save communication and computation resources. Experiment results show that, compared with existing efficient FL frameworks, FLrce improves the computation and communication efficiency by at least 30% and 43% respectively.
@@ 1 Introduction
Even though FL remarkably upholds data privacy, when it comes to application, the effectiveness of FL is usually compromised by the issues of resource shortage and client heterogeneity in real-world IoT networks. On one hand, due to the resource constraints of edge devices, the communication and computation resources of an IoT network are usually finite, presenting a critical challenge to address the resource shortage issue of FL in IoT settings [10, 11, 2]. As shown in Figure 1, first, an IoT network with limited bandwidth support can hardly withstand the exorbitant communication cost in FL when deep-learning models with massive amounts of parameters are communicated frequently between clients and the server [12, 2, 13]. Second, due to the significant energy consumption of training deep learning models on edge devices [14], FL usually generates high computation costs which can be unaffordable for IoT systems with deficient energy supply. To solve the resource shortage problem, an efficient FL framework that can reduce the computation and communication costs without degrading accuracy is expected.
@@
Motivated by the limitation of the state-of-the-art, in this paper, we present FLrce, i.e. Federated Learning with relationship-based client selection and early stopping strategy, to simultaneously address the client heterogeneity and resource shortage problems of FL in IoT settings.
@@
  • We implement FLrce on a physical platform and evaluate FLrce on real-world datasets. The experiment results show that, compared with existing efficient FL frameworks, FLrce at least increases the efficiency of the computation/communication resource utilization by 30% and 43% respectively on average.
@@ 2.1 Federated Learning
Figure 3: Existing efficient FL methods easily become vulnerable to non-iid client data.
(a) Required energy consumption (kJ) per accuracy.
(b) Required data transmission (GB) per accuracy.
Figure 4: Due to the vulnerability to heterogeneous client contributions, existing efficient FL methods consume more communication and computation resources to reach the same target accuracy in the non-iid case.
@@ 2.2 Degraded Performance Caused by Client Heterogeneity
Due to the non-iid local data distribution of FL clients, the performance of existing efficient FL frameworks can be easily compromised by unbalanced training contributions from heterogeneous clients. For illustration, we implement Fedcom [16], Fedprox [21], and Dropout [25] on the CIFAR10 dataset [34], and compare their performances between the cases where the local data distributions are iid and non-iid. In the iid case, all clients share a centralized global dataset. In the non-iid case, data are unevenly distributed to clients following the experimental setting of [35]. The simulation runs on 100 clients with 20 iterations, and other parameter settings can be found in the experiment section. As shown in Figure 3, existing efficient FL methods are significantly impeded by client heterogeneity in the non-iid case. The reason is that, these methods usually employ a trade-off strategy over local clients to reduce the computation or communication cost, such as model pruning [25], training epoch reduction [21] and message compression [16]. While for important clients, applying these strategies might prohibit FL from learning sufficient knowledge from the important local data, and consequently degrades the global model accuracy.
@@ 2.3 Additional Resource Consumption Resulting from Low Learning Efficiency
Owing to the degraded accuracy caused by client heterogeneity, the expected computation cost (energy) and communication cost (data transmission) per accuracy of existing efficient FL methods inevitably escalate as Figure 4 shows. Consequently, given any target accuracy, these methods are expected to consume more computation and communication resources to reach that accuracy. Specifically, Fedcom, Fedprox and Dropout respectively consume 2.45×, 2.09×, 1.75× computation and communication resources to achieve the same target accuracy compared with the iid-case.
@@ 3 Design of FLrce
In this section, we introduce the proposed FLrce framework. Specifically, subsection 3.1 lays out a high-level demonstration of how FLrce works. Subsections 3.2 and 3.3 respectively explain the client selection strategy and the early stopping mechanism of FLrce. Subsection 3.4 summarizes all the details and presents a comprehensive overview of FLrce.
@@ 3.1 Solution Overview
A high-level workflow of FLrce is presented in Figure 5. ①: At the beginning of round $t$, the server decides the set of active clients based on a client selection strategy $h$. $h$ takes $H$ as input and returns the server a set of clients $\mathbb{C}_{t}$, where $H$ is the heuristic map that stores the heuristic values of clients (see Algorithm 2). ②: The server broadcasts the current global model $w^{t}$ to all selected clients. ③: In parallel, each participating client $C_{k}$ ($C_{k}\in\mathbb{C}_{t}$) updates the received parameter by applying stochastic gradient descent (SGD) over its personal dataset (see Equation (3)). ④: Each $C_{k}$ uploads the parameter updates $u^{t}_{k}$ to the server. ⑤: For each $C_{k}\in\mathbb{C}_{t}$, the server writes the corresponding $u^{t}_{k}$ to $V_{k}$, and writes $t$ to $R_{k}$. $V$ is an update map that stores each client’s latest update, $R$ is a time map that stores each client’s last active round. ⑥: The server aggregates the received updates and updates the global model parameter (see Equation (4)). ⑦: The server evaluates the relationship between each selected $C_{k}$ and other clients, and writes the relationships into a relationship map $\Omega\in\mathbb{R}^{M\times M}$, using a relationship modeling function $g$ (see Algorithm 1). ⑧: The server computes each $C_{k}$’s latest heuristic value based on the newest $\Omega$ and updates the heuristic map $H$ (see Equation (7)). ⑨: The server examines whether the early-stopping mechanism has been triggered. If so, the server terminates FL prematurely (see subsection 3.3). If not, the server moves ahead to round $t+1$ and restarts from step ①. The summary of all mathematical notations is listed in Table I.
@@ 3.2 Relationship-based Client Selection
Algorithm 1 Relationship modeling function $g$
1:update $u^{t}_{k}$, old map $\Omega_{k}$, update map $V$, time map $R$, round $t$
@@
6:         synchronous relationship modeling:
@@
9:         asynchronous relationship modeling:
@@
13:return new map $\Omega_{k}$
@@
$d_{p}>d_{o}$ indicates that the global model is possibly moving away from $C_{q}$’s local optimum after incorporating the update from $C_{p}$, then a negative relationship degree will be assigned between $C_{p}$ and $C_{q}$. If $d_{p}<d_{o}$, $C_{p}$ and $C_{q}$ will have a positive relationship degree for a similar reason. To control the relationship degree on the same scale as cosine similarity, we make the result in Equation (6) be no less than -1. The computation of Equation (6) will not infringe on data privacy, as it only requires clients’ parameter updates and the global model parameter. With the synchronous and asynchronous RM combined, the RM function $g$ is implemented in Algorithm 1.
Heuristic Values Computation. The heuristic value of a client $C_{k}$ reflects the importance of $C_{k}$ to the collaborative learning process. Motivated by [15], FLrce evaluates the importance of a client as the combination of its relationships with other peers. Therefore, we define the heuristic of $C_{k}$ as a summation of its relationship degrees with other clients:
$H_{k}=\sum_{j=1,j\neq k}^{M}\Omega_{k,j}$ (7)
@@
3.3 Early Stopping Mechanism
FLrce presents a novel early stopping mechanism settings that terminates FL ahead of the maximum training round. Unlike [33] which uses early stopping to prevent overfitting in centralized machine learning, the early stopping mechanism in FLrce aims to reduce the overall communication and computation resource consumption in decentralized FL settings. The point is that, as FL goes on, the global model tends to ”oscillate” between the important clients’ local optima. In this case, it becomes much more difficult for the global model to reach the global optimum, which triggers the ES mechanism to avoid further consumption of communication and computation resources.
For illustration, Figure 9 shows a simple two-dimensional example. Suppose a network with five clients $C_{1}\sim C_{5}$. $C_{2}$ and $C_{3}$ are the two most important clients as their local optima are closer to the global optimum than others. Given $P=2$, the server tends to to select $C_{2}$ and $C_{3}$ in an exploit round. Suppose two exploit rounds $t_{1}$ and $t_{2}$. At $t_{1}$, the global model $w^{t_{1}}$ is far away from the optimal model $w^{*}$, and the server finds $C_{2}$ and $C_{3}$ not conflicting as $cossim(u_{2}^{t_{1}},u_{3}^{t_{1}})>0$. On the contrary, $w^{t_{2}}$ is closer to $w^{*}$ and the server finds $C_{2}$ and $C_{3}$ conflicting as $cossim(u_{2}^{t_{2}},u_{3}^{t_{2}})<0$. In summary, selecting the important clients can accelerate training in earlier rounds. However, as FL goes on, important clients are likely to conflict with each other and provide inconsistent parameter updates. In this case, selecting the important clients is no longer effective in improving the accuracy of the global model.
@@ 3.3 Early Stopping Mechanism
1:max iterations $T$, number of participants $P$, clients $\mathbb{C}$, explore-exploit factor $\phi$, ES threshold $\psi$
2:Globally, server $S$ executes:
3:initialize $w^{0}$, $H$, $\Omega$, $R$, $V$
4:for $t=1,2,...,T$ do
5:     client selection:
6:           $\mathbb{C}_{t}\leftarrow h(H,T,P,\mathbb{C},\phi)$triangleright Alg. 2
7:     for $k\in\{1,...,M\},C_{k}\in\mathbb{C}_{t}$ do
8:         send $w^{t}$ to $C_{k}$
9:         receive $u^{t}_{k}$ from $C_{k}$
10:         write update/round to the update/time map:
11:               $V_{k}\leftarrow u^{t}_{k},R_{k}\leftarrow t$
12:     end for
13:     aggregation:
14:           $w^{t+1}\leftarrow w^{t}+\sum_{k:C_{k}\in\mathbb{C}_{t}}p_{k}u_{k}^{t}$triangleright Eq. (4)
15:     for $k\in\{1,...,M\},C_{k}\in\mathbb{C}_{t}$ do
16:         update relationship:
17:               $\Omega_{k}\leftarrow g(u^{t}_{k},V,R,t,\Omega_{k})$triangleright Alg. 1
18:         update heuristic:
19:               $H_{k}=\sum_{j=1,j\neq k}^{M}\Omega_{k,j}$triangleright Eq. (7)
20:     end for
21:     check early stopping:
22:     if $ES(t,\mathbb{C}_{t},P,\psi)==$ True then $\triangleright$ Alg. 3
23:         break
24:     end if
25:end for
26:return $w^{t+1}$
@@ 4 Experiment
Figure 10: Mean validation accuracy per round.
Dataset Fedcom Fedprox Dropout PyramidFL TimelyFL FLrce
Accuracy Round Accuracy Round Accuracy Round Accuracy Round Accuracy Round Accuracy Round
EMNIST 78.2% 100 65.9% 100 8.5% 100 81.3% 100 80.06% 100 80.63% 44
Google Speech 55.8% 100 47.1% 100 21.6% 100 64.7% 100 66.8% 100 66.05% 52
CIFAR10 42.9% 100 39.5% 100 14.9% 100 50.19% 100 57.8% 100 57.82% 63
CIFAR100 32.3% 100 29.9% 100 13.7% 100 35.11% 100 36.67% 100 37.57% 47
Average 52.3% 100 45.6% 100 14.6% 100 57.82% 100 60.33% 100 60.51% 52
TABLE III: Total rounds of training and final validation accuracy.
4.1 Experiment Setup
  • Extended MNIST (EMNIST) [38] is an image dataset consisting of 814,255 handwritten characters from 3550 writers. Each sample is a black-and-white-based image with $28\times 28$ pixels, belonging to one of the total 62 categories (10 digits and 52 upper/lower-case letters).
  • CIFAR10 [34] is an image dataset consisting of 50,000 samples covering 10 classes of real-world objects. Each sample is a colorful image with $32\times 32$ pixels.
  • CIFAR100 [34] is an image dataset containing 50,000 samples covering 100 classes of real-world objects. Each sample is a colorful image with $32\times 32$ pixels.
  • Google Speech [39] is a voice dataset consisting of 101,012 samples covering 35 human spoken words. Each sample is a short audio of a word pronounced by one of the total 2618 speakers.
Parameter settings. For EMNIST and Google Speech, we use a convolutional neural network (CNN) with two convolutional layers and one fully-connected layer following [25]. For CIFAR10 and CIFAR100, we use a CNN with two convolutional layers and three fully-connected layers following [27]. The number of global iterations is $T=100$ and the number of clients’ local training epochs is 5 universally following the settings in [23, 16]. The batch size is set to 128 for CIFAR10 and CIFAR100, and 16 for EMNIST and Google Speech by following [26, 23]. We determine the learning rate $\eta$ through hyperparameter tuning. If $\eta$ is too large, all methods reach a high accuracy in the very first few rounds, making it extremely hard to distinguish their differences. If $\eta$ is too small, the training goes slowly and requires much more iterations and time to produce a proper outcome. After hyperparameter tuning, the learning rate $\eta$ is set to 2e-4 for EMNIST, 5e-4 for Google Speech, 0.1 for CIFAR10 and 0.5 for CIFAR100. Lastly, the explore-exploit factor $\phi$ depends on $t$, that is, the explore possibility of FLrce is 1.0 initially and decays by a coefficient of 0.98 every round following [23].
System implementation. The experiment platform is built upon the Flower framework (1.4.0) [40] and Pytorch 2.0.0 [41]. We build a virtual network consisting of one server and $M=100$ clients, with $P=10$ clients participating in FL per round by following the experimental setting in [25]. The server runs on a remote cloud server and selects clients through the interface of Flower. Clients run the local training program on NVIDIA Jetson Nano Developer Kits [42]. We allow the edge devices to switch identities between several clients to conduct large-scale experiments with a limited number of physical devices [26].
Non-iid data distributions. To create a training environment with heterogeneous client contributions, data are unevenly separated into clients following a Dirichlet distribution. Specifically, for EMNIST and Google Speech, data are allocated to clients based on the number of writers/speakers, with each client $k$ accounting for $p_{k}$ of the total 3550/2618 writers or speakers. The distributions of each $p_{k}$ follow a Dirichlet Distribution $Dir(\alpha)$ with a concentration parameter $\alpha$. For CIFAR10 and CIFAR100, the proportions of samples from each class $y$ vary among clients following a Dirichlet Distribution $Dir_{y}(\alpha)$. The parameter $\alpha$ of the Dirichlet Distribution is set to 0.1 following [35].
Baselines. As mentioned before, existing efficient FL methods can be divided into three categories: message compression, accuracy relaxation and dropout. Therefore, we compare FLrce against the typical works in these three categories. Out of resource concerns, we do not include any works that simply address client heterogeneity without taking efficiency into account. The baselines include:
  • Fedcom [16] let clients compress the parameter updates before uploading the updates to the server so as to reduce communication costs.
  • PyramidFL [23] lets clients dynamically reduce local training epochs similar to Fedprox. Additionally, it applies a client selection strategy aiming to accelerate training by selecting clients with larger training losses and less training time.
  • TimelyFL [43] combines local epoch reduction and layer freezing altogether to reduce the computation cost of clients’ local training tasks.
Metrics. We comprehensively compare FLrce with the baselines in terms of accuracy, computation efficiency and communication efficiency. For accuracy, by the end of every global iteration, we test the global model on every local dataset and record the mean validation accuracy. For computation efficiency, we measure the overall energy consumption of the Jetson Nano Develop Kits during the local training process, and we combine the energy consumption with the final accuracy to derive the computation efficiency. For communication efficiency, we measure the bandwidth usage for transmitting the model between clients and the server, and we combine the bandwidth usage with the final accuracy to derive the communication efficiency.
4.2 Experiment Results
4.2.1 Accuracy and Learning efficiency
The experiment outcome shows that FLrce outperforms the baselines in terms of learning efficiency and final accuracy. From Figure 10, we can see that the learning efficiency of FLrce is usually higher than the baselines, representing that FLrce can achieve higher accuracy given the same training rounds. Furthermore, with the help of the ES mechanism, FLrce is able to complete the global training task using only $40\sim 60$% of the maximum global iteration and still achieve equivalent or higher final accuracy. Specifically, as shown by Table III, it takes only 44, 52, 63 and 47 rounds for FLrce to reach an accuracy of 80.06%, 66.8%, 57.8% and 36.67% on EMNIST, Google Speech, CIFAR10 and CIFAR 100 respectively. Comparatively, Fedcom uses 100 rounds to reach an accuracy of 78.2%, 55.8%, 42.9% and 32.3%. Fedprox uses 100 rounds to reach an accuracy of 65.9%, 47.1%, 39.5% and 29.9%. Dropout uses 100 rounds to reach an accuracy of 8.5%, 21.6%, 14.9% and 13.7%, which are completely exceeded by FLrce. Moreover, even though PyramidFL and TimelyFL achieve higher final accuracy than FLrce occasionally, they require much more training rounds and therefore more resources to reach that accuracy. Overall, even with fewer training rounds, FLrce still achieves the highest accuracy CIFAR10 and CIFAR100 and the second highest accuracy on EMNIST and Google Speech. On average, FLrce achieves the highest final accuracy with at least 0.18% improvement compared with the baselines.
(a) EMNIST
(b) Google Speech
(c) CIFAR10
(d) CIFAR100